# Deep Inverse Design for High-Level Synthesis

### v1

@@ Abstract
High-level synthesis (HLS) has significantly advanced the automation of digital circuits design, yet the need for expertise and time in pragma tuning remains challenging. Existing solutions for the design space exploration (DSE) adopt either heuristic methods, lacking essential information for further optimization potential, or predictive models, missing sufficient generalization due to the time-consuming nature of HLS and the exponential growth of the design space. To address these challenges, we propose Deep Inverse Design for HLS (DID4444HLS), a novel approach that integrates graph neural networks and generative models. DID4HLS iteratively optimizes hardware designs aimed at compute-intensive algorithms by learning conditional distributions of design features from post-HLS data. Compared to four state-of-the-art DSE baselines, our method achieved an average improvement of $42.5\%$ on average distance to reference set (ADRS) compared to the best-performing baselines across six benchmarks, while demonstrating high robustness and efficiency.
@@ I Introduction
The increasing complexity of circuit designs, compounded by stringent time-to-market imperatives, underscores the demand for advanced design automation [1]. High-level synthesis (HLS) is an increasingly popular and pivotal solution that facilitates the conversion of algorithms expressed in high-level languages (HLLs) such as C/C++/SystemC into hardware descriptions [2]. However, HLS as a multi-objective optimization problem (MOOP) necessitates expertise and experience to navigate the optimization among the conflicting objectives such as latency and area [3, 4]. Consequently, there is a high demand for the automation of the design space exploration (DSE) process to address the complex tuning challenges for pragmas (or directives).
Heuristic methods offer rapid implementations to yield improved designs [5, 6, 7, 8, 9, 10]. However, their effectiveness is limited by the lack of a thorough understanding of the behaviors of HLS tools. We need to understand how operations are scheduled in order to estimate design performance, but the scheduling information is only available after the HLS process [11]. Recently, graph neural networks (GNNs) have been employed for code analysis to predict synthesis results, demonstrating advantages over the preceding DSE methods [12, 13, 14, 15, 16]. Compilers transform algorithms in HLL into platform-independent intermediate representations (IRs) that encapsulate semantic information [17, 18, 19]. They are subsequently transformed into control and data flow graphs (CDFGs) for GNN training [20, 21]. However, three critical issues hinder the effectiveness of these methods:
@@
$2$) Current methods are tied to specific IR compilers. Using these methods in a different toolchain introduces representation discrepancies. For example, GNNs adopting LLVM IR [17] does not match HLS tools operating Hector [23]. Besides, even when the compilers match, changes in compiler versions require additional workload on the methods. These discrepancies lead to ineffective graph feature extraction [3].
@@
To address these challenges, we propose a novel method called Deep Inverse Design for HLS (DID4444HLS). DID4444HLS focuses on choosing the best pragma configurations of the compute-intensive algorithms to achieve designs that minimize the objectives of latency and area. We construct the design space with the pragmas of pipelining, loop unrolling, array partitioning, and function inlining. The uniqueness of DID4444HLS lies in its strategy to work on a distribution that samples pragma configurations, with constraints that prevent redundant design samples according to the strategies of HLS tools, such as interactions between pipelining and loop unrolling. Initially, DID4444HLS uses a Uniform pragma distribution to explore a wide range of design possibilities. It then iteratively synthesizes batches of designs generated from the updated pragma distributions, resulting in the IRs that are generally represented in CDFG and crucial in deriving their final hardware designs [28, 29]. In this work, we call these IRs as post-HLS IRs (PIRs). These PIRs, along with data on their objectives, are collected into a dataset which is used to train the objective predictors based on Graph Attention Network v2222 (GATv2222) [30]. Furthermore, DID4444HLS adopts Variational Autoencoder (VAE) to learn the distribution of design features that determine design performance. They are extracted from the PIRs by the GATv2222 models. This training is conditioned on the pragma distributions. Through this process, DID4444HLS effectively converges the pragma distribution to approximate Pareto front [3].
@@
$1$) By realizing inverse design, DID4444HLS requires only a limited synthesis budget to achieve well-optimized designs, thereby significantly speeding up the synthesis process.
@@
We evaluated DID4444HLS through rigorous experiments and quantified its performance in comparison to several prior works. Our results revealed significant improvements in the quality of the Pareto fronts generated by DID4444HLS. In comparison to the best-performing DSE baselines, DID4444HLS improves the average distance to the reference set (ADRS) by $42.5\%$ on average over the baselines for six benchmarks adapted from Polybench [31] and MachSuite [32]. That is, DID4444HLS significantly outperforms the state-of-the-art in achieving a close Pareto front to the optimal. The comparative experiments also demonstrate the high robustness and efficiency of our method. Furthermore, to validate the construction of DID4444HLS, we also conducted an ablation study to explore the benefits of using PIRs and VAE. Results show that our approach achieved an average improvement of $48.0\%$ over the best-performing ablation configurations.
@@ III Problem Definition
Because HLS optimization is a MOOP, a single distribution is unable to cover the whole Pareto front. For instance, in the context of the trade-off between latency and area, the distribution prioritizing low latency might be significantly different from the one prioritizing low area. In order to approximate Pareto front, we pre-define a set of optimization weights $\bm{\lambda}$ sweeping from 00 to $1$ and we denote the $k$-$th$ weight as $\bm{\lambda}_{k}$.
@@
$\bm{c}_{k}=\bm{\lambda}_{k}l+(1-\bm{\lambda}_{k})a$ (4)
where $l$ and $a$ denote the standardized latency and area of $d$. The mean and standard deviation of latency and area for their standardization are obtained from the designs that are high-level synthesized during the experiments.
We create a vector set $\bm{\Theta}$ in which the vector $\bm{\Theta}_{k}$ corresponds to $\bm{\lambda}_{k}$, as a pragma distribution from which pragma configurations are sampled. Each configuraion combines with the algorithm in HLL into a design $d$ from a pre-defined design space $S$. Our goal is to identify $\bm{\Theta}_{k}$ that yields designs close to the optimal design:
@@ IV-A Design Space Interface
To start the optimization process targeting pipelining, loop unrolling, array partitioning, and function inlining, an interface is established between algorithms in C++ and our method. Loops within the C++ code files are labeled. The arrays, denoted by $\mathbf{arr}$, introduced below are user-defined that determine the design space $S$. $\mathbf{arr}_{nest}$ describes how the loops nest. An example code with loop labels for HLS is shown in Fig. 2. $\mathbf{arr}_{nest}=[-1,0,1]$ means the structure of the three loops. Here, $\mathbf{arr}_{nest}[2]=1$, $\mathbf{arr}_{nest}[1]=0$, and $\mathbf{arr}_{nest}[0]=-1$ mean that the loop of index $2$ is nested under the loop of index $1$, and both are nested under the loop of index 00. $\mathbf{arr}_{unroll}$ contains user-defined loop unrolling factors. For example, $\mathbf{arr}_{unroll}=[[1,2],[1,2,4],[1,2,4,8]]$ means that $\mathbf{arr}_{unroll}[0]=[1,2]$ allows the loop of index 00 to be unrolled with a factor of $1$ or $2$, and the other two subarrays are for the rest loops in order. $\mathbf{arr}_{pipeline}$ contains probabilities for each loop to be pipelined. $\mathbf{arr}_{ii}$ contains initiation intervals for a loop if it is pipelined. $\mathbf{arr}_{interface}$ specifies the arrays, at the interface of the top functions, to be partitioned with the type of block or cyclic. $\mathbf{arr}_{p\_itf}$ contains their array partitioning factors. $\mathbf{arr}_{inline}$ contains probabilities for each function to be inlined. $S$ determines the initialization of $\bm{\Theta}$ which is then mapped to align with the deep models, rendering the method inherently self-adaptive. To maximize exploration, the probabilities within $\bm{\Theta}$ to sample the pragma values are initialized to Uniform at the beginning of the DID4444HLS process. New designs are synthesized to augment the dataset $D$ in each iteration. To avoid excessive workloads from redundant design samples with $\mathbf{arr}_{bound}$ containing the loop bounds, we constrain the design sampling process under three HLS rules [39, 33]:
@@ IV-B DID4HLS
The pseudocode in Alg. 1 summarizes DID4444HLS. A set of optimization weights $\bm{\lambda}$ sweeping from 00 to $1$ is pre-defined in order to approximate Pareto front. The arrays introduced in the section IV-A are pre-defined to determine the design space $S$ for each benchmark, and then a set of distributions $\bm{\Theta}$, each $\bm{\Theta}_{k}$ corresponding to $\bm{\lambda}_{k}$, are initialized to Uniform (line $3$). $n_{1}$ pragma configurations are sampled from $\bm{\Theta}$ under the constraint of the three sampling rules. The algorithm in C++ code combining each configuration gives a design $d$. All of them form a set $\{d\}_{n_{1}}$ (line $6$). We invoke the HLS tool to synthesize the designs from $\{d\}_{n_{1}}$ that are new to $D$ (lines $7$-$8$). The synthesis results include PIR CDFG $g$, design latency $l$ and area $a$, and the $\bm{\Theta}$ vectors under which the configurations are sampled (line $11$). They are collected into $D$ (line $12$). Then, $l$ and $a$ are standardized (line $13$). Two predictors $P_{l}(g)$ and $P_{a}(g)$ are trained with $D$ for predicting standardized $l$ and $a$, respectively (line $14$). The model training takes $g$ as input, standardized $l$ and $a$ as ground truth. Each predictor comprises two parts: an extractor, $F_{l}(g)$ or $F_{a}(g)$, for extracting graph feature $\mathbf{f}_{l}$ or $\mathbf{f}_{a}$, and an MLP, $MLP_{l}(\mathbf{f}_{l})$ or $MLP_{a}(\mathbf{f}_{a})$, for predicting $l$ or $a$. After training the predictors, $\mathbf{f}_{l}$ and $\mathbf{f}_{a}$ of the synthesized designs are extracted (line $15$), combined with corresponding $\bm{\Theta}$ vectors, for training two estimators $V_{l}$ and $V_{a}$ that reconstruct $\mathbf{f}_{l}$ and $\mathbf{f}_{a}$ conditioned on $\bm{\Theta}_{k}$ (line $16$). Then, we build a surrogate based on $MLP_{l}$, $MLP_{a}$, and the decoders $Dec_{l}$ of $V_{l}$ and $Dec_{a}$ of $V_{a}$. $n_{2}$ prior samples $\{\mathbf{z}_{i}\}$, combined with $\bm{\Theta}_{k}$ of current iteration, result in $\hat{\bm{c}}_{k}$ as the estimation of the cost $\bm{c}_{k}$ (lines $18$-$21$). Through gradient descent with respect to $\hat{\bm{c}}_{k}$, $\bm{\Theta}_{k}$ is updated for several times, where we build a function $\eta$ with the aid of the ADAM optimizer [40] to output updating rates while guaranteeing validity of $\bm{\Theta}_{k}$ (line $22$). In the next iteration, if the synthesis budget $B$ is reached, an approximate Pareto front will be returned (lines $9$-$10$).
@@
14:    Standardize $l,a\in D$
15:    Train $P_{l}$ and $P_{a}$ with $g,l,a\in D$
16:    Extract $\mathbf{f}_{l}=F_{l}(g)$ and $\mathbf{f}_{a}=F_{a}(g)$ with $g\in D$
17:    Train $V_{l}$ and $V_{a}$ with $\mathbf{f}_{l}$, $\mathbf{f}_{a}$, and [$\bm{\Theta}_{k}]_{d}\in D$
@@
21:        $\hat{a}_{i}=MLP_{a}(Dec_{a}(\mathbf{z}_{i},\bm{\Theta}_{k}))$
22:        $\hat{\bm{c}}_{ki}=\bm{\lambda}_{k}\hat{l}_{i}+(1-\bm{\lambda}_{k})\hat{a}_{i}$
@@ IV-C Predictors
Fig. 4 illustrates the predictor architecture. We construct $F_{l}(g)$ and $F_{a}(g)$, each with three layers of GATv2222 which demonstrates state-of-the-art performance when applied to graph data [30]. An ELU activation layer is applied between two adjacent GATv2222 layers [41]. Before the PIR data are input into the deep models, their node and edge attributes are embedded.
@@ IV-D Estimators
Following the training of $P_{l}$ and $P_{a}$, $F_{l}$ and $F_{a}$ extract graph features $\mathbf{f}$ of $g\in D$. Then $\mathbf{f}$ and $\bm{\Theta}$ vectors under which the designs are sampled are adopted to train $V_{l}$ and $V_{a}$ based on VAE for learning conditional distributions of $\mathbf{f}$. The architecture of these estimators is illustrated in Fig. 5, wherein $Dec$ realizes $\mu$ in Eq. 3, $\mathbf{m}$ and $\mathbf{v}$ realize $m_{i}$ and $v_{i}$ respectively in Eq. 3 across all dimensions $i$. We assume the prior distribution as $p(\mathbf{z})=N(\mathbf{0},\mathbf{I})$ [37]. After training, the combination of a prior sample $\mathbf{z}\sim p(\mathbf{z})$ and $\bm{\Theta}_{k}$ transformed by an MLP projector as input into a decoder yields a sample of $\mathbf{f}$ from the estimated distribution.

### v2

@@ Abstract
High-level synthesis (HLS) has significantly advanced the automation of digital circuits design, yet the need for expertise and time in pragma tuning remains challenging. Existing solutions for the design space exploration (DSE) adopt either heuristic methods, lacking essential information for further optimization potential, or predictive models, missing sufficient generalization due to the time-consuming nature of HLS and the exponential growth of the design space. To address these challenges, we propose Deep Inverse Design for HLS (DID4444HLS), a novel approach that integrates graph neural networks and generative models. DID4HLS iteratively optimizes hardware designs aimed at compute-intensive algorithms by learning conditional distributions of design features from post-HLS data. Compared to four state-of-the-art DSE baselines, our method achieved an average improvement of $42.8\%$ on average distance to reference set (ADRS) compared to the best-performing baselines across six benchmarks, while demonstrating high robustness and efficiency.
@@ I Introduction
The increasing complexity of circuit designs, compounded by stringent time-to-market imperatives, underscores the demand for advanced design automation [1]. High-level synthesis (HLS) is an increasingly popular and pivotal solution that facilitates the conversion of algorithms expressed in high-level languages (HLLs) such as C/C++/SystemC into hardware descriptions [2]. However, HLS as a multi-objective optimization problem (MOOP) necessitates expertise and experience to navigate the optimization among the conflicting objectives such as latency and resource consumption [3, 4]. Consequently, there is a high demand for the automation of the design space exploration (DSE) process to address the complex tuning challenges for pragmas (or directives).
Heuristic methods offer rapid implementations to yield improved designs [5, 6, 7, 8, 9, 10]. However, their effectiveness is limited by the lack of a thorough understanding of the behaviors of HLS tools. We need to understand how operations are scheduled in order to estimate design performance, but scheduling information is only available after the HLS process [11]. Recently, graph neural networks (GNNs) have been employed for code analysis to predict synthesis results, demonstrating advantages over the preceding DSE methods [12, 13, 14, 15, 16]. Compilers transform algorithms in HLL into platform-independent intermediate representations (IRs) that encapsulate semantic information [17, 18, 19]. They are subsequently transformed into control and data flow graphs (CDFGs) for GNN training [20, 21]. However, three critical issues hinder the effectiveness of these methods:
@@
$2$) Current methods are tied to specific IR compilers. Using these methods in a different toolchain introduces representation discrepancies. For example, GNNs adopting LLVM IR [17] does not match HLS tools operating MLIR [23]. Besides, even when the compilers match, changes in compiler versions require additional workload on the methods. These discrepancies lead to ineffective graph feature extraction [3].
@@
To address these challenges, we propose a novel method called Deep Inverse Design for HLS (DID4444HLS). DID4444HLS focuses on choosing the optimal pragma configurations of the compute-intensive algorithms to achieve designs that are close to Pareto points. We construct the design space with the pragmas of pipelining, loop unrolling, array partitioning, and function inlining. The uniqueness of DID4444HLS lies in its strategy to work on a distribution that samples pragma configurations, with constraints that prevent redundant design samples according to the strategies of HLS tools, such as interactions between pipelining and loop unrolling. Initially, DID4444HLS uses a random pragma distribution to explore a wide range of design possibilities. It then iteratively synthesizes batches of designs generated from the updated pragma distributions, resulting in the IRs that are generally represented in CDFG and crucial in deriving their final hardware designs [28, 29]. In this work, we call these IRs as post-HLS IRs (PIRs). These PIRs, along with data on their objectives, are collected into a dataset which is used to train the objective predictors based on Graph Attention Network v2222 (GATv2222) [30]. Furthermore, DID4444HLS adopts Variational Autoencoder (VAE) to learn the distribution of design features that determine design performance. They are extracted from the PIRs by the GATv2222 models. This training is conditioned on the pragma distributions. Through this process, DID4444HLS effectively converges the pragma distribution to approximate Pareto front [3].
@@
$1$) By realizing inverse design, DID4444HLS requires a modest synthesis budget — the limitation of the number of synthesized designs throughout the optimization process — to achieve well-optimized designs, thereby significantly speeding up the synthesis process.
@@
We evaluated DID4444HLS through rigorous experiments and quantified its performance in comparison to several prior works. Our results revealed significant improvements in the quality of the Pareto fronts generated by DID4444HLS. In comparison to the best-performing DSE baselines, DID4444HLS improves the average distance to the reference set (ADRS) by $42.8\%$ on average for six benchmarks adapted from Polybench [31] and MachSuite [32]. That is, DID4444HLS significantly outperforms the state-of-the-art in achieving a close Pareto front to the optimal. The comparative experiments also demonstrate the high robustness and efficiency of our method. Furthermore, to validate the construction of DID4444HLS, we also conducted an ablation study to explore the benefits of using PIRs and VAE. Results show that our approach achieved an average improvement of $43.0\%$ over the best-performing ablation configurations.
@@ III Problem Definition
Because HLS optimization is a MOOP, a single distribution is unable to cover the whole Pareto front. For instance, in the context of the trade-off between latency and resource consumption, the distribution prioritizing low latency might be significantly different from the one prioritizing low resource consumption. In order to approximate Pareto front, we pre-define a set of optimization weights $\bm{\lambda}$ sweeping from 00 to $1$ and we denote the $k$-$th$ weight as $\bm{\lambda}_{k}$.
@@
$\bm{c}_{k}=\bm{\lambda}_{k}l+(1-\bm{\lambda}_{k})r$ (4)
where $l$ and $r$ denote the standardized latency and resource consumption of $d$. The mean and standard deviation of latency and resource consumption for their standardization are obtained from the designs that are high-level synthesized during the experiments.
We create a vector set $\bm{\Theta}$ in which the vector $\bm{\Theta}_{k}$ corresponds to $\bm{\lambda}_{k}$, as a pragma distribution from which pragma configurations are sampled. Each configuration combines with the algorithm in HLL into a design $d$ from a pre-defined design space $S$. Our goal is to identify $\bm{\Theta}_{k}$ that yields designs close to the optimal design:
@@ IV-A Design Space Interface
To start the optimization process targeting pipelining, loop unrolling, array partitioning, and function inlining, an interface is established between algorithms in C++ and our method. Loops within the C++ code files are labeled. The arrays, denoted by $\mathbf{arr}$, introduced below are user-defined that determine the design space $S$. $\mathbf{arr}_{nest}$ describes how the loops nest. An example code with loop labels for HLS is shown in Fig. 2. $\mathbf{arr}_{nest}=[-1,0,1]$ means the structure of the three loops. Here, $\mathbf{arr}_{nest}[2]=1$, $\mathbf{arr}_{nest}[1]=0$, and $\mathbf{arr}_{nest}[0]=-1$ mean that the loop of index $2$ is nested under the loop of index $1$, and both are nested under the loop of index 00. $\mathbf{arr}_{unroll}$ contains user-defined loop unrolling factors. For example, $\mathbf{arr}_{unroll}=[[1,2],[1,2,4],[1,2,4,8]]$ means that $\mathbf{arr}_{unroll}[0]=[1,2]$ allows the loop of index 00 to be unrolled with a factor of $1$ or $2$, and the other two subarrays are for the rest loops in order. $\mathbf{arr}_{ii}$ contains initiation intervals for a loop if it is pipelined. $\mathbf{arr}_{interface}$ specifies the arrays, at the interface of the top functions, to be partitioned with the type of block or cyclic. $\mathbf{arr}_{p\_itf}$ contains their array partitioning factors. $\mathbf{arr}_{inline}$ contains probabilities for each function to be inlined. $S$ determines the initialization of $\bm{\Theta}$ which is then mapped to align with the deep models, rendering the method inherently self-adaptive. To avoid excessive workloads from redundant design samples with $\mathbf{arr}_{bound}$ containing the loop bounds, we constrain the design sampling process under three HLS rules [39, 33]:
@@
To maximize exploration, the probabilities within $\bm{\Theta}$ to sample the pragma values are initialized to Uniform at the beginning of the DID4444HLS process for fully random sampling, except for $\mathbf{arr}_{pipeline}$ that contains probabilities for each loop to be pipelined. The resource required increases exponentially with the depth of loop nests when unrolling outmost loops. For example, if we unroll loop_$\_0$ in Fig. 2, all its inner loops will be fully unrolled. This often results in a challenge where HLS tools are unable to complete the synthesis process in limited time. Therefore, we initialize $\mathbf{arr}_{pipeline}$ with low probabilities in general, and the outer loops are unrolled with lower probabilities. After sampling, new designs are synthesized to augment the dataset $D$ in each iteration.
@@ IV-B DID4HLS
The pseudocode in Alg. 1 summarizes DID4444HLS. A set of optimization weights $\bm{\lambda}$ sweeping from 00 to $1$ is pre-defined in order to approximate Pareto front. The arrays introduced in the section IV-A are pre-defined to determine the design space $S$ for each benchmark, and then a set of distributions $\bm{\Theta}$, each $\bm{\Theta}_{k}$ corresponding to $\bm{\lambda}_{k}$, are initialized to Uniform (line $3$). $n_{1}$ pragma configurations are sampled from $\bm{\Theta}$ under the constraint of the three sampling rules. The algorithm in C++ code combining each configuration gives a design $d$. All of them form a set $\{d\}_{n_{1}}$ (line $6$). We invoke the HLS tool to synthesize the designs from $\{d\}_{n_{1}}$ that are new to $D$ (lines $7$-$8$). The synthesis results include PIR CDFG $g$, design latency $l$ and resource consumption $r$, and the $\bm{\Theta}$ vectors under which the configurations are sampled (line $11$). They are collected into $D$ (line $12$). Then, $l$ and $r$ are standardized (line $13$). Two predictors $P_{l}(g)$ and $P_{r}(g)$ are trained with $D$ for predicting standardized $l$ and $r$, respectively (line $14$). The model training takes $g$ as input, standardized $l$ and $r$ as ground truth. Each predictor comprises two parts: an extractor, $F_{l}(g)$ or $F_{r}(g)$, for extracting graph feature $\mathbf{f}_{l}$ or $\mathbf{f}_{r}$, and an MLP, $MLP_{l}(\mathbf{f}_{l})$ or $MLP_{r}(\mathbf{f}_{r})$, for predicting $l$ or $r$. After training the predictors, $\mathbf{f}_{l}$ and $\mathbf{f}_{r}$ of the synthesized designs are extracted (line $15$), combined with corresponding $\bm{\Theta}$ vectors, for training two estimators $V_{l}$ and $V_{r}$ that reconstruct $\mathbf{f}_{l}$ and $\mathbf{f}_{r}$ conditioned on $\bm{\Theta}_{k}$ (line $16$). Then, we build a surrogate based on $MLP_{l}$, $MLP_{r}$, and the decoders $Dec_{l}$ of $V_{l}$ and $Dec_{r}$ of $V_{r}$. $n_{2}$ prior samples $\{\mathbf{z}_{i}\}$, combined with $\bm{\Theta}_{k}$ of current iteration, result in $\hat{\bm{c}}_{k}$ as the estimation of the cost $\bm{c}_{k}$ (lines $18$-$21$). Through gradient descent with respect to $\hat{\bm{c}}_{k}$, $\bm{\Theta}_{k}$ is updated for several times, where we build a function $\eta$ with the aid of the ADAM optimizer [40] to output updating rates while guaranteeing validity of $\bm{\Theta}_{k}$ (line $22$). In the next iteration, if the synthesis budget $B$ is reached, an approximate Pareto front will be returned (lines $9$-$10$).
@@
14:    Standardize $l,r\in D$
15:    Train $P_{l}$ and $P_{r}$ with $g,l,r\in D$
16:    Extract $\mathbf{f}_{l}=F_{l}(g)$ and $\mathbf{f}_{r}=F_{r}(g)$ with $g\in D$
17:    Train $V_{l}$ and $V_{r}$ with $\mathbf{f}_{l}$, $\mathbf{f}_{r}$, and [$\bm{\Theta}_{k}]_{d}\in D$
@@
21:        $\hat{r}_{i}=MLP_{r}(Dec_{r}(\mathbf{z}_{i},\bm{\Theta}_{k}))$
22:        $\hat{\bm{c}}_{ki}=\bm{\lambda}_{k}\hat{l}_{i}+(1-\bm{\lambda}_{k})\hat{r}_{i}$
@@ IV-C Predictors
Fig. 4 illustrates the predictor architecture. We construct $F_{l}(g)$ and $F_{r}(g)$, each with three layers of GATv2222 which demonstrates state-of-the-art performance when applied to graph data [30]. An ELU activation layer is applied between two adjacent GATv2222 layers [41]. Before the PIR data are input into the deep models, their node and edge attributes are embedded.
@@ IV-D Estimators
Following the training of $P_{l}$ and $P_{r}$, $F_{l}$ and $F_{r}$ extract graph features $\mathbf{f}$ of $g\in D$. Then $\mathbf{f}$ and $\bm{\Theta}$ vectors under which the designs are sampled are adopted to train $V_{l}$ and $V_{r}$ based on VAE for learning conditional distributions of $\mathbf{f}$. The architecture of these estimators is illustrated in Fig. 5, wherein $Dec$ realizes $\mu$ in Eq. 3, $\mathbf{m}$ and $\mathbf{v}$ realize $m_{i}$ and $v_{i}$ respectively in Eq. 3 across all dimensions $i$. We assume the prior distribution as $p(\mathbf{z})=N(\mathbf{0},\mathbf{I})$ [37]. After training, the combination of a prior sample $\mathbf{z}\sim p(\mathbf{z})$ and $\bm{\Theta}_{k}$ transformed by an MLP projector as input into a decoder yields a sample of $\mathbf{f}$ from the estimated distribution.